### v2

# Deep Learning for Camera Calibration and Beyond: A Survey

Abstract
Camera calibration involves estimating camera parameters to infer geometric features from captured sequences, which is crucial for computer vision and robotics. However, conventional calibration is laborious and requires dedicated collection. Recent efforts show that learning-based solutions have the potential to be used in place of the repeatability works of manual calibrations. Among these solutions, various learning strategies, networks, geometric priors, and datasets have been investigated. In this paper, we provide a comprehensive survey of learning-based camera calibration techniques, by analyzing their strengths and limitations. Our main calibration categories include the standard pinhole camera model, distortion camera model, cross-view model, and cross-sensor model, following the research trend and extended applications. As there is no benchmark in this community, we collect a holistic calibration dataset that can serve as a public platform to evaluate the generalization of existing methods. It comprises both synthetic and real-world data, with images and videos captured by different cameras in diverse scenes. Toward the end of this paper, we discuss the challenges and provide further research directions. To our knowledge, this is the first survey for the learning-based camera calibration (spanned 8 years). The summarized methods, datasets, and benchmarks are available and will be regularly updated at https://github.com/KangLiao929/Awesome-Deep-Camera-Calibration.
Index Terms:
Camera calibration, Deep learning, Computational photography, Multiple view geometry, 3D vision, Robotics.1 Introduction
Camera calibration is a fundamental and indispensable field in computer vision and it has a long research history [1, 2, 3, 4], tracing back to around 60 years ago[5]. The first step for many vision and robotics tasks is to calibrate the intrinsic (image sensor and distortion parameters) and/or extrinsic (rotation and translation) camera parameters, ranging from computational photography, and multi-view geometry, to 3D reconstruction. In terms of the task type, there are different techniques to calibrate the standard pinhole camera, fisheye lens camera, stereo camera, light field camera, event camera, and LiDAR-camera system, etc. Figure 1 shows the popular calibration objectives, models, and extended applications in camera calibration.
Traditional methods for camera calibration generally depend on hand-crafted features and model assumptions. These methods can be broadly divided into three categories. The most prevalent one involves using a known calibration target (e.g., a checkerboard) as it is deliberately moved in the 3D scene [6, 7, 8]. Then, the camera captures the target from different viewpoints and the checkerboard corners are detected for calculating the camera parameters. However, such a procedure requires cumbersome manual interactions and it cannot achieve automatic calibration “in the wild”. To pursue better flexibility, the second category of camera calibration, i.e., the geometric-prior-based calibration has been largely studied [9, 10, 11, 12]. To be specific, the geometric structures are leveraged to model the 3D-2D correspondence in the scene, such as lines and vanishing points. However, this type of method heavily relies on structured man-made scenes containing rich geometric priors, leading to poor performance when applied to general environments. The third category is self-calibration[13, 14, 15]. Such a solution takes a sequence of images as inputs and estimates the camera parameters using multi-view geometry. The accuracy of self-calibration, however, is constrained by the limits of the feature detectors, which can be influenced by diverse lighting conditions and textures. The above parametric models are based on the physical interpretation of camera geometry. While they are user-friendly, they tend to be tailored for specific camera models and might not offer optimal accuracy. Instead, non-parametric models [16, 17, 18] link each image pixel to its 3D observation ray, eliminating the constraints of parametric models.
Since there are many standard techniques for calibrating cameras in an industry/laboratory implementation[19, 20], this process is usually ignored in recent developments. However, calibrating single and wild images remains challenging, especially when images are collected from websites and unknown camera models. This challenge motivates the researchers to investigate a new paradigm.
Recently, deep learning has brought new inspirations to camera calibration and its applications. Learning-based methods achieve state-of-the-art performances on various tasks with higher efficiency. In particular, diverse deep neural networks (DNNs) have been developed, such as convolutional neural networks (CNNs), generative adversarial networks (GANs), PointNet, and vision transformers (ViTs), of which the high-level semantic features show more powerful representation capability compared with the hand-crafted features. Moreover, diverse learning strategies have been exploited to boost the geometric perception of neural networks. Learning-based methods offer a flexible and end-to-end camera calibration solution, without manual interventions or calibration targets, which sets them apart from traditional methods. Furthermore, some of these methods achieve camera model-free and label-free calibration, showing promising and meaningful applications.
With the rapid increase in the number of learning-based camera calibration methods, it has become increasingly challenging to keep up with new advances. Consequently, there is an urgent need to analyze existing works and foster a community dedicated to this field. Previously, some surveys, e.g., [21, 22, 23] only focused on a specific task/camera in camera calibration or one type of approach. For instance, Salvi et al. [21] reviewed the traditional camera calibration methods in terms of the algorithms. Hughes et al. [22] provided a detailed review for calibrating fisheye cameras with traditional solutions. While Fan et al. [23] discussed both the traditional methods and deep learning methods, their survey only considers calibrating the wide-angle cameras. In addition, due to the few amount of reviewed learning-based methods (around 10 papers), the readers are difficult to picture the development trend of general camera calibration in Fan et al. [23].
In this paper, we provide a comprehensive and in-depth overview of recent advances in learning-based camera calibration, covering over 100 papers. We also discuss potential directions for further improvements and examine various types of cameras and targets. To facilitate future research on different topics, we categorize the current solutions according to calibration objectives and applications. In addition to fundamental parameters such as focal length, rotation, and translation, we also provide detailed reviews for correcting image distortion (radial distortion and rolling shutter distortion), estimating cross-view mapping, calibrating camera-LiDAR systems, and other applications. Such a trend follows the development of cameras and market demands for virtual reality, autonomous driving, neural rendering, etc.
To our best knowledge, this is the first survey of the learning-based camera calibration and its extended applications, it has the following unique contributions. (1) Our work mainly follows recent advances in deep learning-based camera calibration. In-depth analysis and discussion in various aspects are offered, including publications, network architecture, loss functions, datasets, evaluation metrics, learning strategies, implementation platforms, etc. The detailed information of each literature is listed in Table I. (2) Despite the calibration algorithm, we comprehensively review the classical camera models and their extended models. In particular, we summarize the redesigned calibration objectives in deep learning since some traditional calibration objectives are verified to be hard to learn by neural networks. (3) We collect a dataset containing images and videos captured by different cameras in different environments, which can serve as a platform to evaluate the generalization of existing methods. (4) We discuss the open challenges of learning-based camera calibration and propose some future directions to provide guidance for further research in this field. (5) An open-source repository is created that provides a taxonomy of all reviewed works and benchmarks. The repository will be updated regularly in https://github.com/KangLiao929/Awesome-Deep-Camera-Calibration.
In the following sections, we discuss and analyze various aspects of learning-based camera calibration. The remainder of this paper is organized as follows. In Section 2, we provide the concrete learning paradigms and learning strategies of the learning-based camera calibration. Subsequently, we introduce and discuss the specific methods based on the standard camera model, distortion model, cross-view model, and cross-sensor model in Section 3, Section 4, Section 5, and Section 6, respectively (see Figure 2). The collected benchmark for calibration methods is depicted in Section 7. Finally, we conclude the learning-based camera calibration and suggest the future directions of this community in Section 8.
2 Preliminaries
Deep learning has brought new inspirations to camera calibration, enabling a fully automatic calibration procedure without manual intervention. Here, we first summarize two prevalent paradigms in learning-based camera calibration: regression-based calibration and reconstruction-based calibration. Then, the widely-used learning strategies are reviewed in this research field. The detailed definitions for classical camera models and their corresponding calibration objectives are exhibited in the supplementary material.
2.1 Learning Paradigm
Driven by different architectures of the neural network, the researchers have developed two main paradigms for learning-based camera calibration and its applications.
Regression-based Calibration Given an uncalibrated input, the regression-based calibration first extracts the high-level semantic features using stacked convolutional layers. Then, the fully connected layers aggregate the semantic features and form a vector of the estimated calibration objective. The regressed parameters are used to conduct subsequent tasks such as distortion rectification, image warping, camera localization, etc. This paradigm is the earliest and has a dominant role in learning-based camera calibration and its applications. All the first works in various objectives, e.g., intrinsics: Deepfocal [24], extrinsic: PoseNet [25], radial distortion: Rong et al. [26], rolling shutter distortion: URS-CNN [27], homography matrix: DHN [28], hybrid parameters: Hold-Geoffroy et al. [29], camera-LiDAR parameters: RegNet [30] have been achieved with this paradigm.
Reconstruction-based Calibration On the other hand, the reconstruction-based calibration paradigm discards the parameter regression and directly learns the pixel-level mapping function between the uncalibrated input and target, inspired by the conditional image-to-image translation [31] and dense visual perception[32, 33]. The reconstructed results are then calculated for the pixel-wise loss with the ground truth. In this regard, most reconstruction-based calibration methods [34, 35, 36, 37] design their network architecture based on the fully convolutional network such as U-Net[38]. Specifically, an encoder-decoder network, with skip connections between the encoder and decoder features at the same spatial resolution, progressively extracts the features from low-level to high-level and effectively integrates multi-scale features. At the last convolutional layer, the learned features are aggregated into the target channel, reconstructing the calibrated result at the pixel level.
In contrast to the regression-based paradigm, the reconstruction-based paradigm does not require the label of diverse camera parameters. Besides, the imbalance loss problem can be eliminated since it only optimizes the photometric loss of calibrated results. Therefore, the reconstruction-based paradigm enables a blind camera calibration without a strong camera model assumption.
2.2 Learning Strategies
In the following, we review the learning-based camera calibration literature regarding different learning strategies.
Supervised Learning Most learning-based camera calibration methods train their networks with the supervised learning strategy, from the classical methods [24, 25, 28, 39, 26, 40] to the state-of-the-art methods [41, 42, 43, 44, 45]. In terms of the learning paradigm, this strategy supervises the network with the ground truth of the camera parameters (regression-based paradigm) or paired data (reconstruction-based paradigm). In general, they synthesize the training dataset from other large-scale datasets, under the random parameter/transformation sampling and camera model simulation. Some recent works [46, 47, 48, 49] establish their training dataset using a real-world setup and label the captured images with manual annotations, thereby fostering advancements in this research domain.
Semi-Supervised Learning Training the network using an annotated dataset under diverse scenarios is an effective learning strategy. However, human annotation can be prone to errors, leading to inconsistent annotation quality or the inclusion of contaminated data. Consequently, increasing the training dataset to improve performance can be challenging due to the complexity and cost of constructing the dataset. To address this challenge, SS-WPC[50] proposes a semi-supervised method for correcting portraits captured by a wide-angle camera. It employs a surrogate task (segmentation) and a semi-supervised method that utilizes direction and range consistency and regression consistency to leverage both labeled and unlabeled data.
Weakly-Supervised Learning Although significant progress has been made, data labeling for camera calibration is a notorious costly process, and obtaining perfect ground-truth labels is challenging. As a result, it is often preferable to use weak supervision with machine learning methods. Weakly supervised learning refers to the process of building prediction models through learning with inadequate supervision. Zhu et al. [51] present a weakly supervised camera calibration method for single-view metrology in unconstrained environments, where there is only one accessible image of a scene composed of objects of uncertain sizes. This work leverages 2D object annotations from large-scale datasets, where people and buildings are frequently present and serve as useful “reference objects” for determining 3D size.
Unsupervised Learning Unsupervised learning, commonly referred to as unsupervised machine learning, analyzes and groups unlabeled datasets using machine learning algorithms. UDHN [52] is the first work for a cross-view camera model using unsupervised learning, which estimates the homography matrix of a paired image without the projection labels. By reducing a pixel-wise intensity error that does not require ground truth data, UDHN [52] outperforms previous supervised learning techniques. While preserving superior accuracy and robustness to fluctuation in light, the proposed unsupervised algorithm can also achieve faster inference time. Inspired by this work, increasing more methods leverage the unsupervised learning strategy to estimate the homography such as CA-UDHN [53], BaseHomo [54], HomoGAN[55], and Liu et al. [56]. Besides, UnFishCor [57] frees the demands for distortion parameters and designs an unsupervised framework for the wide-angle camera.
Self-supervised Learning Robotics is where the phrase “self-supervised learning” first appears, as training data is automatically categorized by utilizing relationships between various input sensor signals. Compared to supervised learning, self-supervised learning leverages input data itself as the supervision. Many self-supervised techniques are presented to learn visual characteristics from massive amounts of unlabeled photos or videos without the need for time-consuming and expensive human annotations. SSR-Net [58] presents a self-supervised deep homography estimation network, which relaxes the need for ground truth annotations and leverages the invertibility constraints of homography. To be specific, SSR-Net [58] utilizes the homography matrix representation in place of other approaches’ typically-used 4-point parameterization, to apply the invertibility constraints. SIR [59] devises a brand-new self-supervised camera calibration pipeline for wide-angle image rectification, based on the principle that the corrected results of distorted images of the same scene taken with various lenses need to be the same. With self-supervised depth and pose learning as a proxy aim, Fang et al. [60] present to self-calibrate a range of generic camera models from raw video, offering for the first time a calibration evaluation of camera model parameters learned solely via self-supervision.
Reinforcement Learning Instead of aiming to minimize at each stage, reinforcement learning can maximize the cumulative benefits of a learning process as a whole. To date, DQN-RecNet [61] is the first and only work in camera calibration using reinforcement learning. It applies a deep reinforcement learning technique to tackle the fisheye image rectification by a single Markov Decision Process, which is a multi-step gradual calibration scheme. In this situation, the current fisheye image represents the state of the environment. The agent, Deep Q-Network [62], generates an action that should be executed to correct the distorted image.
In the following, we will review the specific methods and literature for learning-based camera calibration. The structural and hierarchical taxonomy is shown in Figure 2.
Method | Publication | Objective | Network | Loss Function | Dataset | Evaluation | Learning | Platform | Simulation | |
2015 | DeepFocal [24] | ICIP | Standard | AlexNet | loss | 1DSfM[63] | Accuracy | SL | Caffe | |
PoseNet [25] | ICCV | Standard | GoogLeNet | loss | Cambridge Landmarks[64] | Accuracy | SL | Caffe | ||
2016 | DeepHorizon [65] | BMVC | Standard | GoogLeNet | Huber loss | HLW[66] | Accuracy | SL | Caffe | |
DeepVP [39] | CVPR | Standard | AlexNet | Logistic loss | YUD[67], ECD[68], HLW[66] | Accuracy | SL | Caffe | ||
Rong et al. [26] | ACCV | Distortion | AlexNet | Softmax loss | ImageNet[69] | Line length | SL | Caffe | ✓ | |
DHN[28] | RSSW | Cross-View | VGG | loss | MS-COCO[70] | MSE | SL | Caffe | ✓ | |
2017 | CLKN [71] | CVPR | Cross-View | CNNs | Hinge loss | MS-COCO[70] | MSE | SL | Torch | ✓ |
HierarchicalNet [72] | ICCVW | Cross-View | VGG | loss | MS-COCO[70] | MSE | SL | TensorFlow | ✓ | |
URS-CNN [27] | CVPR | Distortion | CNNs | loss | Sun[73], Oxford[74], Zubud[75], LFW[76] | PSNR, RMSE | SL | Torch | ✓ | |
RegNet [30] | IV | Cross-Sensor | CNNs | loss | KITTI[77] | MAE | SL | Caffe | ✓ | |
2018 | Hold-Geoffroy et al. [29] | CVPR | Standard | DenseNet | Entropy loss | SUN360[78] | Human sensitivity | SL | - | |
DeepCalib [40] | CVMP | Distortion | Inception-V3 | Logcosh loss | SUN360[78] | Mean error | SL | TensorFlow | ✓ | |
FishEyeRecNet [79] | ECCV | Distortion | VGG | loss | ADE20K[80] | PSNR, SSIM | SL | Caffe | ✓ | |
Shi et al.[81] | ICPR | Distortion | ResNet | loss | ImageNet[69] | MSE | SL | PyTorch | ✓ | |
DeepFM[82] | ECCV | Cross-View | ResNet | loss | T&T[83], KITTI[77], 1DSfM[63] | F-score, Mean | SL | PyTorch | ✓ | |
Poursaeed et al.[84] | ECCVW | Cross-View | CNNs | , loss | KITTI[77] | EPI-ABS, EPI-SQR | SL | - | ||
UDHN[52] | RAL | Cross-View | VGG | loss | MS-COCO[70] | RMSE | USL | TensorFlow | ✓ | |
PFNet[85] | ACCV | Cross-View | FCN | Smooth loss | MS-COCO[70] | MAE | SL | TensorFlow | ✓ | |
CalibNet[86] | IROS | Cross-Sensor | ResNet | Point cloud distance, loss | KITTI[77] | Geodesic distance, MAE | SL | TensorFlow | ✓ | |
Chang et al.[87] | ICRA | Standard | AlexNet | Cross-entropy loss | DeepVP-1M [87] | MSE, Accuracy | SL | Matconvnet | ||
2019 | Lopez et al. [88] | CVPR | Distortion | DenseNet | Bearing loss | SUN360[78] | MSE | SL | PyTorch | |
UprightNet [89] | ICCV | Standard | U-Net | Geometry loss | InteriorNet[90], ScanNet[91], SUN360[78] | Mean error | SL | PyTorch | ||
Zhuang et al. [92] | IROS | Distortion | ResNet | loss | KITTI[77] | Mean error, RMSE | SL | PyTorch | ✓ | |
SSR-Net [58] | PRL | Cross-View | ResNet | loss | MS-COCO[70] | MAE | SSL | PyTorch | ✓ | |
Abbas et al. [93] | ICCVW | Cross-View | CNNs | Softmax loss | CARLA[94] | AUC[95], Mean error | SL | TensorFlow | ✓ | |
DR-GAN [34] | TCSVT | Distortion | GANs | Perceptual loss | MS-COCO[70] | PSNR, SSIM | SL | TensorFlow | ✓ | |
STD [96] | TCSVT | Distortion | GANs+CNNs | Perceptual loss | MS-COCO[70] | PSNR, SSIM | SL | TensorFlow | ✓ | |
Deep360Up [97] | VR | Standard | DenseNet | Log-cosh loss[98] | SUN360[78] | Mean error | SL | - | ✓ | |
UnFishCor [57] | JVCIR | Distortion | VGG | loss | Places2[99] | PSNR, SSIM | USL | TensorFlow | ✓ | |
BlindCor [37] | CVPR | Distortion | U-Net | loss | Places2[99] | MSE | SL | PyTorch | ✓ | |
RSC-Net [100] | CVPR | Distortion | ResNet | loss | KITTI[77] | Mean error | SL | PyTorch | ✓ | |
Xue et al. [101] | CVPR | Distortion | ResNet | loss | Wireframes[102], SUNCG[103] | PSNR, SSIM, RPE | SL | PyTorch | ✓ | |
Zhao et al. [46] | ICCV | Distortion | VGG+U-Net | loss | Self-constructed+BU-4DFE[104] | Mean error | SL | - | ✓ | |
NeurVPS [105] | NeurIPS | Standard | CNNs | Binary cross entropy, chamfer- loss | ScanNet [91], SU3 [106] | Angle accuracy | SL | PyTorch | ||
2020 | Sha et al. [107] | CVPR | Cross-View | U-Net | Cross-entropy loss | World Cup 2014[108] | IoU | SL | TensorFlow | |
Lee et al. [109] | ECCV | Standard | PointNet + CNNs | Cross-entropy loss | Google Street View[110], HLW[66] | Mean error, AUC[95] | SL | - | ||
MisCaliDet [111] | ICRA | Distortion | CNNs | loss | KITTI[77] | MSE | SL | TensorFlow | ✓ | |
DeepPTZ [112] | WACV | Distortion | Inception-V3 | loss | SUN360[78] | Mean error | SL | PyTorch | ✓ | |
MHN [113] | CVPR | Cross-View | VGG | Cross-entropy loss | MS-COCO[70], Self-constructed | MAE | SL | TensorFlow | ✓ | |
Davidson et al. [114] | ECCV | Standard | FCN | Dice loss | SUN360[78] | Accuracy | SL | - | ✓ | |
CA-UDHN [53] | ECCV | Cross-View | FCN + ResNet | Triplet loss | Self-constructed | MSE | USL | PyTorch | ||
DeepFEPE [115] | IROS | Standard | VGG + PointNet | loss | KITTI[77], ApolloScape[116] | Mean error | SL | PyTorch | ||
DDM [35] | TIP | Distortion | GANs | loss | MS-COCO[70] | PSNR, SSIM | SL | TensorFlow | ✓ | |
Li et al. [117] | TIP | Distortion | CNNs | Cross-entropy, loss | CelebA[118] | Cosine distance | SL | - | ✓ | |
PSE-GAN [119] | ICPR | Distortion | GANs | , WGAN loss | Place2[99] | MSE | SL | - | ✓ | |
RDC-Net [120] | ICIP | Distortion | ResNet | , loss | ImageNet[69] | PSNR, SSIM | SL | PyTorch | ✓ | |
FE-GAN [121] | ICASSP | Distortion | GANs | , GAN loss | Wireframe[102], LSUN[122] | PSNR, SSIM, RMSE | SSL | PyTorch | ✓ | |
RDCFace [123] | CVPR | Distortion | ResNet | Cross-entropy, loss | IMDB-Face[124] | Accuracy | SL | - | ✓ | |
LaRecNet [125] | arXiv | Distortion | ResNet | loss | Wireframes[102], SUNCG[103] | PSNR, SSIM, RPE | SL | PyTorch | ✓ | |
Baradad et al. [126] | CVPR | Standard | CNNs | loss | ScanNet[91], NYU[127], SUN360[78] | Mean error, RMS | SL | PyTorch | ||
Zheng et al. [128] | CVPR | Standard | CNNs | loss | FocaLens[129] | Mean error, PSNR, SSIM | SL | - | ✓ | |
Zhu et al. [51] | ECCV | Standard | CNNs + PointNet | loss | SUN360[78], MS-COCO[70] | Mean error, Accuracy | WSL | PyTorch | ✓ | |
DeepUnrollNet [49] | CVPR | Distortion | FCN | , perceptual, total variation loss | Carla-RS[49], Fastec-RS[49] | PSNR, SSIM | SL | PyTorch | ✓ | |
RGGNet [130] | RAL | Cross-Sensor | ResNet | Geodesic distance loss | KITTI[77] | MSE, MSEE, MRR | SL | TensorFlow | ✓ | |
CalibRCNN [131] | IROS | Cross-Sensor | RNNs | , Epipolar geometry loss | KITTI [77] | MAE | SL | TensorFlow | ✓ | |
SSI-Calib [132] | ICRA | Cross-Sensor | CNNs | loss | Pascal VOC 2012 [133] | Mean/standard deviation | SL | TensorFlow | ✓ | |
SOIC [134] | arXiv | Cross-Sensor | ResNet + PointRCNN | Cost function | KITTI [77] | Mean error | SL | - | ||
NetCalib [135] | ICPR | Cross-Sensor | CNNs | loss | KITTI [77] | MAE | SL | PyTorch | ✓ | |
SRHEN [136] | ACM-MM | Cross-View | CNNs | loss | MS-COCO [70], SUN397 [78] | MACE | SL | - | ✓ | |
2021 | StereoCaliNet [137] | TCI | Standard | U-Net | loss | TAUAgent[138], KITTI[77] | Mean error | SL | PyTorch | ✓ |
CTRL-C [139] | ICCV | Standard | Transformer | Cross-entropy, loss | Google Street View[110], SUN360[78] | Mean error, AUC[95] | SL | PyTorch | ✓ | |
Wakai et al. [140] | ICCVW | Distortion | DenseNet | Smooth loss | StreetLearn[141] | Mean error, PSNR, SSIM | SL | - | ✓ | |
OrdianlDistortion [142] | TIP | Distortion | CNNs | Smooth loss | MS-COCO[70] | PSNR, SSIM, MDLD | SL | TensorFlow | ✓ | |
PolarRecNet [143] | TCSVT | Distortion | VGG + U-Net | , loss | MS-COCO[70], LMS[144] | PSNR, SSIM, MSE | SL | PyTorch | ✓ | |
DQN-RecNet [61] | PRL | Distortion | VGG | loss | Wireframes[102] | PSNR, SSIM, MSE | RL | PyTorch | ✓ | |
Tan et al. [47] | CVPR | Distortion | U-Net | loss | Self-constructed | Accuracy | SL | PyTorch | ||
PCN [145] | CVPR | Distortion | U-Net | , , GAN loss | Place2[99] | PSNR, SSIM, FID, CW-SSIM | SL | PyTorch | ✓ | |
DaRecNet [36] | ICCV | Distortion | U-Net | Smooth , loss | ADE20K[80] | PSNR, SSIM | SL | PyTorch | ✓ | |
DLKFM [146] | CVPR | Cross-View | Siamese-Net | loss | MS-COCO[70], Google Earth, Google Map | MSE | SL | TensorFlow | ✓ | |
LocalTrans [147] | ICCV | Cross-View | Transformer | loss | MS-COCO[70] | MSE, PSNR, SSIM | SL | PyTorch | ✓ | |
BasesHomo [54] | ICCV | Cross-View | ResNet | Triplet loss | CA-UDHN[53] | MSE | USL | PyTorch | ||
ShuffleHomoNet [148] | ICIP | Cross-View | ShuffleNet | loss | MS-COCO[70] | RMSE | SL | TensorFlow | ✓ | |
DAMG-Homo [44] | TCSVT | Cross-View | CNNs | loss | MS-COCO[70], UDIS[149] | RMSE, PSNR, SSIM | SL | TensorFlow | ✓ | |
SA-MobileNet [150] | BMVC | Standard | MobileNet | Cross-entropy loss | SUN360[78], ADE20K[80], NYU[127] | MAE, Accuracy | SL | TensorFlow | ✓ | |
SPEC [48] | ICCV | Standard | ResNet | Softargmax- loss | Self-constructed | W-MPJPE, PA-MPJPE | SL | PyTorch | ✓ | |
DirectionNet [151] | CVPR | Standard | U-Net | Cosine similarity loss | InteriorNet[90], Matterport3D[152] | Mean and median error | SL | TensorFlow | ✓ | |
JCD [153] | CVPR | Distortion | FCN | Charbonnier[154], perceptual loss | BS-RSCD [153], Fastec-RS [49] | PSNR, SSIM, LPIPS | SL | PyTorch | ||
LCCNet [155] | CVPRW | Cross-Sensor | CNNs | Smooth , loss | KITTI[77] | MSE | SL | PyTorch | ✓ | |
CFNet [156] | Sensors | Cross-Sensor | FCN | , Charbonnier[154] loss | KITTI[77], KITTI-360[157] | MAE, MSEE, MRR | SL | PyTorch | ✓ | |
Fan et al. [158] | ICCV | Distortion | U-Net | , perceptual loss | Carla-RS [49], Fastec-RS [49] | PSNR, SSIM, LPIPS | SL | PyTorch | ||
SUNet [159] | ICCV | Distortion | DenseNet + ResNet | , perceptual loss | Carla-RS [49], Fastec-RS [49] | PSNR, SSIM | SL | PyTorch | ||
SemAlign [160] | IROS | Cross-Sensor | CNNs | Semantic alignment loss | KITTI [77] | Mean/median rotation errors | SL | PyTorch | ✓ | |
2022 | DVPD [41] | CVPR | Standard | CNNs | Cross-entropy loss | SU3[106], ScanNet[91], YUD[67], NYU[127] | Accuracy, AUC[95] | SL | PyTorch | ✓ |
Fang et al. [60] | ICRA | Standard | CNNs | loss | KITTI[77], EuRoC[161], OmniCam[162] | MRE, RMSE | SSL | PyTorch | ||
CPL [163] | ICASSP | Standard | Inception-V3 | loss | CARLA[94], CyclistDetection[164] | MAE | SL | TensorFlow | ✓ | |
IHN [165] | CVPR | Cross-View | Siamese-Net | loss | MS-COCO[70], Google Earth, Google Map | MACE | SL | PyTorch | ✓ | |
HomoGAN [55] | CVPR | Cross-View | GANs | Cross-entropy, WGAN loss | CA-UDHN[53] | Mean error | USL | PyTorch | ✓ | |
SS-WPC [50] | CVPR | Distortion | Transformer | Cross-entropy, loss | Tan et al.[47] | Accuracy | Semi-SL | PyTorch | ||
AW-RSC [166] | CVPR | Distortion | CNNs | Charbonnier[154], perceptual loss | Self-constructed, FastecRS[49] | PSNR, SSIM | SL | PyTorch | ||
EvUnroll [42] | CVPR | Distortion | U-Net | Charbonnier, perceptual, TV loss | Self-constructed, FastecRS[49] | PSNR, SSIM, LPIPS | SL | PyTorch | ||
Do et al. [167] | CVPR | Standard | ResNet | , Robust angular [168] loss | Self-constructed, 7-SCENES[169] | Median error, Recall | SL | PyTorch | ||
DiffPoseNet [170] | CVPR | Standard | CNNs + LSTM | loss | TartanAir[171], KITTI[77], TUM-RGBD[172] | PEE, AEE[173] | SSL | PyTorch | ||
SceneSqueezer [174] | CVPR | Standard | Transformer | loss | RobotCar Seasons[175], Cambridge Landmarks[64] | Mean error, Recall[173] | SL | PyTorch | ||
FocalPose [176] | CVPR | Standard | CNNs | , Huber loss | Pix3D[177], CompCars[178], StanfordCars[178] | Median error, Accuracy | SL | PyTorch | ||
DXQ-Net [179] | arXiv | Cross-Sensor | CNNs + RNNs | , geodesic loss | KITTI[77], KITTI-360[157] | MSE | SL | PyTorch | ✓ | |
SST-Calib [45] | ITSC | Cross-Sensor | CNNs | loss | KITTI[77] | QAD, AEAD | SL | PyTorch | ✓ | |
CCS-Net [180] | IROS | Distortion | U-Net | loss | TUM-RGBD[172] | MAE, RPE | SL | PyTorch | ✓ | |
FishFormer [43] | arXiv | Distortion | Transformer | loss | Place2[99], CelebA[118] | PSNR, SSIM, FID | SL | PyTorch | ✓ | |
SIR [59] | TIP | Distortion | ResNet | loss | ADE20K[80], WireFrames[102], MS-COCO[70] | PSNR, SSIM | SSL | PyTorch | ✓ | |
ATOP [181] | TIV | Cross-Sensor | CNNs | Cross entropy loss | Self-constructed + KITTI[77] | RRE, RTE | SL | - | ||
FusionNet [182] | ICRA | Cross-Sensor | CNNs+PointNet | loss | KITTI[77] | MAE | SL | PyTorch | ✓ | |
RGKCNet [183] | TIM | Cross-Sensor | CNNs+PointNet | loss | KITTI[77] | MSE | SL | PyTorch | ✓ | |
GenCaliNet [184] | ECCV | Distortion | DenseNet | loss | StreetLearn[141], SP360[185] | MAE, PSNR, SSIM | SL | - | ✓ | |
Liu et al. [56] | TPAMI | Cross-View | ResNet | Triplet loss | Self-constructed | MSE, Accuracy | USL | PyTorch |
3 Standard Model
Generally, for learning-based calibration works, the objectives of the intrinsics calibration contain focal length and optical center, and the objectives of the extrinsic calibration contain the rotation matrix and translation vector.
3.1 Intrinsics Calibration
Deepfocal [24] is a pioneer work in learning-based camera calibration, it aims to estimate the focal length of any image “in the wild”. In detail, Deepfocal considered a simple pinhole camera model and regressed the horizontal field of view using a deep convolutional neural network. Given the width of an image, the relationship between the horizontal field of view and focal length can be described by:
Due to component wear, temperature fluctuations, or outside disturbances like collisions, the calibrated parameters of a camera are susceptible to change over time. To this end, MisCaliDet [111] proposed to identify if a camera needs to be recalibrated intrinsically. Compared to the conventional intrinsic parameters such as the focal length and image center, MisCaliDet presented a new scalar metric, i.e., the average pixel position difference (APPD) to measure the degree of camera miscalibration, which describes the mean value of the pixel position differences over the entire image.
3.2 Extrinsics Calibration
In contrast to intrinsic calibration, extrinsic calibration infers the spatial correspondence of the camera and its located 3D scene. PoseNet[25] first proposed deep convolutional neural networks to regress 6-DoF camera pose in real-time. A pose vector p was predicted by PoseNet, given by the 3D position x and orientation represented by quaternion q of a camera, namely, . For constructing the training dataset, the labels are automatically calculated from a video of the scenario using a structure from motion method [186].
Inspired by PoseNet[25], the following works improved the extrinsic calibration in terms of the intermediate representation, interpretability, data format, learning objective, etc. For example, to optimize the geometric pose objective, DeepFEPE [115] designed an end-to-end keypoint-based framework with learnable modules for detection, feature extraction, matching, and outlier rejection. Such a pipeline imitated the traditional baseline, in which the final performance can be analyzed and improved by the intermediate differentiable module. To bridge the domain gap between the extrinsic objective and image features, recent works proposed to first learn an intermediate representation from the input, such as surface geometry [89], depth map [137], directional probability distribution [151], and normal flow [170], etc. Then, the extrinsic are reasoned by geometric constraints and learned representation. Therefore, the neural networks are gradually guided to perceive the geometry-related features, which are crucial for extrinsic estimation. Considering the privacy concerns and limited storage problem, some recent works compressed the scene and exploited the point-like feature to estimate the extrinsic. For example, Do et al. [167] trained a network to recognize sparse but significant 3D points, dubbed scene landmarks, by encoding their appearance as implicit features. And the camera pose can be calculated using a robust minimal solver followed by a Levenberg-Marquardt-based nonlinear refinement. SceneSqueezer [174] compressed the scene information from three levels: the database frames are clustered using pairwise co-visibility information, a point selection module prunes each cluster based on estimation performance, and learned quantization further compresses the selected points.
3.3 Joint Intrinsic and Extrinsic Calibration
3.3.1 Geometric Representations
Vanishing Points The intersection of projections of a set of parallel lines in the world leads to a vanishing point. The detection of vanishing points is a fundamental and crucial challenge in 3D vision. In general, vanishing points reveal the direction of 3D lines, allowing the agent to deduce 3D scene information from a single 2D image.
DeepVP [39] is the first learning-based work for detecting the vanishing points given a single image. It reversed the conventional process by scoring the horizon line candidates according to the vanishing points they contain. Chang et al. [87] redesigned this task as a CNN classification problem using an output layer with 225 discrete possible vanishing point locations. For constructing the dataset, the camera view is panned and tilted with step 5° from -35° to 35° in the panorama scene (total 225 images) from a single GPS location. To directly leverage the geometric properties of vanishing points, NeurVPS [105] proposed a canonical conic space and a conic convolution operator that can be implemented as regular convolutions in this space, where the learning model is capable of calculating the global geometric information of vanishing points locally. To overcome the need for a large amount of training data in previous methods, DVPD [41] incorporated the neural network with two geometric priors: Hough transformation and Gaussian sphere. First, the convolutional features are transformed into a Hough domain, mapping lines to distinct bins. The projection of the Hough bins is then extended to the Gaussian sphere, where lines are transformed into great circles and vanishing points are located at the intersection of these circles. Geometric priors are data-efficient because they eliminate the necessity for learning this information from data, which enables an interpretable learning framework and generalizes better to domains with slightly different data distributions.
Horizon Lines The horizon line is a crucial contextual attribute for various computer vision tasks especially image metrology, computational photography, and 3D scene understanding. The projection of the line at infinity onto any plane that is perpendicular to the local gravity vector determines the location of the horizon line.
Given the FoV, pitch, and roll of a camera, it is straightforward to locate the horizon line in its captured image space. DeepHorizon [65] proposed the first learning-based solution for estimating the horizon line from an image, without requiring any explicit geometric constraints or other cues. To train the network, a new benchmark dataset, Horizon Lines in the Wild (HLW), was constructed, which consists of real-world images with labeled horizon lines. SA-MobileNet [150] proposed an image tilt detection and correction with self-attention MobileNet [187] for smartphones. A spatial self-attention module was devised to learn long-range dependencies and global context within the input images. To address the difficulty of the regression task, they trained the network to estimate multiple angles within a narrow interval of the ground truth tilt, penalizing only those values that locate outside this narrow range.
3.3.2 Composite Parameters
Calibrating the composite parameters aims to estimate the intrinsic parameters and extrinsic parameters simultaneously. By jointly estimating composite parameters and training using data from a large-scale panorama dataset [78], Hold-Geoffroy et al. [29] largely outperformed previous independent calibration tasks. Moreover, Hold-Geoffroy et al. [29] performed human perception research in which the participants were asked to evaluate the realism of 3D objects composited with and without accurate calibration. This data was further designed to a new perceptual measure for the calibration errors. In terms of the feature category, Lee et al. [109] and CTRL-C [139] considered both semantic features and geometric cues for camera calibration. They showed how taking use of geometric features, is capable of facilitating the network to comprehend the underlying perspective structure of an image. The pipeline of CTRL-C is illustrated in Figure 3. In recent literature, more applications are jointly studied with camera calibration, for example, single view metrology [51], 3D human pose and shape estimation [48], depth estimation [126, 60], object pose estimation [176], and image reflection removal [128], etc.
Considering the heterogeneousness and visual implicitness of different camera parameters, CPL [163] estimated the parameters using a novel camera projection loss, exploiting the camera model neural network to reconstruct the 3D point cloud. The proposed loss addressed the training imbalance problem by representing different errors of camera parameters in terms of a unified metric.
3.4 Discussion
3.4.1 Technique Summary
The above methods target automatic calibration without manual intervention and scene assumption. Early literature [24, 25] separately studied the intrinsic calibration or extrinsic calibration. Driven by large-scale datasets and powerful networks, subsequent works [39, 65, 29, 139] considered a comprehensive camera calibration, inferring various parameters and geometric representations. To relieve the difficulty of learning the camera parameters, some works [89, 137, 151, 170] proposed to learn an intermediate representation. In recent literature, more applications are jointly studied with camera calibration [51, 48, 126, 60, 128]. This suggests solving the downstream vision tasks, especially in 3D tasks may require prior knowledge of the image formation model. Moreover, some geometric priors [41] can alleviate the data-starved requirement of deep learning, showing the potential to bridge the gap between the calibration target and semantic features.
It is interesting to find that increasing more extrinsic calibration methods [115, 167, 174] revisited and restored the traditional feature point-based solutions. The standard extrinsics that describe the camera motion contain limited degrees of freedom, and thus some local features can well represent the spatial correspondence. Besides, the network designed for point learning significantly improves the efficiency of calibration models, such as PointNet [188] and PointCNN [189]. Such a pipeline also enables clear interpretability of learning-based camera calibration, which promotes understanding of how the network calibrates and magnifies the influences of intermediate modules.
NeRF, recognized for its groundbreaking capability in synthesizing novel views from 2D images, has seen various advancements recently. Progress in this arena includes the incorporation of additional trainable components, prior constraints, enhanced network designs, and novel training strategies. Specifically, NeRF [190] marked a pivotal moment, demonstrating the simultaneous optimization of camera parameters and poses during training. It introduced a trainable pinhole camera model, followed by SCNeRF [191] which proposed a more complex camera model, comprising of a pinhole design, radial distortion, and a pixel-specific noise model. While increased trainable parameters offer enhanced representational capabilities, they also complicate training. To counteract this, researchers have utilized prior geometric knowledge as constraints to stabilize optimization, integrating depth estimation [192, 193], multi-view correspondence [191, 194, 193], and GAN-based constraints [195]. Additionally, refinements to NeRF’s architecture have emerged, leveraging Gaussian [196] or sinusoidal activations [197]. Concurrently, training strategies such as coarse-to-fine pipelines [198, 192] and advanced sampling techniques [198, 199, 197] have been proposed, enhancing both reconstruction quality and parameter precision.
Popular frameworks like InstantNGP [200] and NeRFStudio [201] offer features to fine-tune camera parameters. Typically, NeRF methods use outputs from tools like COLMAP or Polycam, with rendering quality tied closely to initial data quality. Minor errors in intrinsics or poses can lead to noticeable rendering artifacts. Current approaches [200, 201] integrate intrinsic parameters, distortion coefficients, and camera poses into NeRF’s training optimization, enhancing both geometric structure learning and parameter optimization. This boosts 3D geometry robustness and refines camera models. However, fine-tuning parameters may complicate optimization and can destabilize training. It may also not work well if there are not enough views to reliably adjust the intrinsics without overfitting.
3.4.2 Future Effort
(1) Explore more model priors. Most learning-based calibration methods study the parametric camera models but their generalization abilities are limited. In contrast, non-parametric models directly model the relationship between the 3D imaging ray and its resulting pixel in the image, encoding valuable priors in the learned semantic features to reason the camera parameters. Recent works [202, 203, 204] also incorporate the perspective of modeling pixel-wise information for camera calibration, making minimal assumptions on the camera model and showing more interpretable and in line with how humans perceive.
(2) Decouple different stages in an end-to-end calibration learning model. Most learning-based camera calibration methods include a feature extraction stage and an objective estimation stage. However, how the networks learn the features related to calibration is ambiguous. Therefore, decoupling the learning process by different traditional calibration stages can guide the way of feature extraction. It would be meaningful to extend the idea in extrinsic calibration [115, 167, 174] to more general calibration problems.
(3) Transfer the measurement space from the parameter error to the geometric difference. When it comes to jointly calibrating various camera parameters, the training process will suffer from an imbalance loss optimization problem. The main reason is different camera parameters correspond to different sample distributions. The simple normalization strategy cannot unify their error spaces. Therefore, we can formulate a straightforward measurement space in terms of the geometric properties of different camera parameters.
(4) Despite the strides achieved by recent NeRF methods, training NeRF without precise camera parameters remains a challenge, especially in scenarios with sparse views, pronounced movements, low-texture regions, and suboptimal initial values. Contemporary NeRF-based methodologies do optimize camera parameters, yielding notable results. However, they exhibit significant computational demands and lack the generalization seen in current deep-learning calibration techniques. We contend that in the present NeRF-based methods, camera parameters often play a secondary role. Thus, crafting effective calibration algorithms that capitalize on NeRF remains an arduous but promising endeavor.
4 Distortion Model
In the learning-based camera calibration, calibrating the radial distortion and roll shutter distortion gains increasing attention due to their widely used applications for the wide-angle lens and CMOS sensor. In this part, we mainly review the calibration/rectification of these two distortions.
4.1 Radial Distortion
The literature on learning-based radial distortion calibration can be classified into two main categories: regression-based solutions and reconstruction-based solutions.
4.1.1 Regression-based Solution
Rong et al. [26] and DeepCalib [40] are pioneer works for the learning-based wide-angle camera calibration. They treated the camera calibration as a supervised classification [26] or regression [40] problem, and then the networks with the convolutional layers and fully connected layers were used to learn the distortion features of inputs and predict the camera parameters. In particular, DeepCalib [40] explored three learning solutions for wide-angle camera calibration as illustrated in Figure 4. Their experiments showed the simplest architecture SingleNet achieves the best performance on both accuracy and efficiency. To enhance the distortion perception of networks, the following works investigated introducing more diverse features such as the semantic features [79] and geometry features [101, 125, 123]. Additionally, some works improved the generalization by designing learning strategies such as unsupervised learning [57], self-supervised learning [59], and reinforcement learning [46]. By randomly chosen coefficients throughout each mini-batch of the training process, RDC-Net [120] was able to dynamically generate distortion images on-the-fly. It enhanced the rectification performance and prevents the learning model from overfitting. Instead of contributing to the techniques of deep learning, other works leaned to explore the vision prior to interpretable calibration. For example, having observed the radial distortion image owns the center symmetry characteristics, in which the texture far from the image center has stronger distortion, Shi et al. [81] and PSE-GAN [119] developed a position-aware weight layer (fixed [81] and learnable [119]) of this property and enabled the network to explicitly perceive the distortion. Lopez et al. [88] proposed a novel parameterization for radial distortion that is better suited for networks than directly learning the distortion parameters. Furthermore, OrdinalDistortion [142] presented a learning-friendly representation, i.e., ordinal distortion. Compared to the implicit and heterogeneous camera parameters, such a representation can facilitate the distortion perception of the neural network due to its clear relation to the image features.
4.1.2 Reconstruction-based Solution
Inspired by the conditional image-to-image translation and dense visual perception, the reconstruction-based solution starts to evolve from the conventional regression-based paradigm. DR-GAN [34] is the first reconstruction-based solution for calibrating the radial distortion, which directly models the pixel-wise mapping between the distorted image and rectified image. It achieved the camera parameter-free training and one-stage rectification. Thanks to the liberation of the assumption of camera models, the reconstruction-based solution showed the potential to calibrate various types of cameras in one learning network. For example, DDM [35] unified different camera models into a domain by presenting the distortion distribution map, which explicitly describes the distortion level of each pixel in a distorted image. Then, the network learned to reconstruct the rectified image using this geometric prior map. To make the mapping function interpretable, the subsequent works [96, 37, 46, 121, 145, 47, 50, 143] developed the displacement filed between the distorted image and rectified image. Such a manner is able to eliminate the generated artifacts in the pixel-level reconstruction. In particular, FE-GAN [121] integrated the geometry prior like Shi et al. [81] and PSE-GAN [119] into their reconstruction-based solution and presented a self-supervised strategy to learn the distortion flow for wide-angle camera calibration in Figure 5. Most reconstruction-based solutions exploit a U-Net-like architecture to learn pixel-level mapping. However, the distortion feature can be transferred from encoder to decoder by the skip-connection operation, leading to a blurring appearance and incomplete correction in reconstruction results. To address this issue, Li et al. [117] abandoned the skip-connection in their rectification network. To keep the feature fusion and restrain the geometric difference simultaneously, PCN [145] designed a correction layer in skip-connection and applied the appearance flows to revise the convolved features in different encoder layers. Having noticed that the previous sampling strategy of the convolution kernel neglected the radial symmetry of distortion, PolarRecNet [143] transformed the distorted image from the Cartesian coordinates domain into the polar coordinates domain.
4.2 Roll Shutter Distortion
The existing deep learning calibration works on roll shutter (RS) distortion can be classified into two categories: single-frame-based [27, 100, 42] and multi-frame-based [49, 153, 159, 158, 166]. The single-frame-based solution studies the case of a single roll shutter image as input and directly learns to correct the distortion using neural networks. The ideal corrected result can be regarded as the global shutter (GS) image. It is an ill-posed problem and requires some additional prior assumptions to be defined. On the contrary, the multi-frame-based solution considers the consecutive frames (two or more) of a video taken by a roll shutter camera, in which the strong temporal correlation can be investigated for more reasonable correction.
4.2.1 Single-frame-based Solution
URS-CNN [27] is the first learning work for calibrating the rolling shutter camera. In this work, a neural network with long kernel characteristics was used to understand how the scene structure and row-wise camera motion interact. To specifically address the nature of the RS effect produced by the row-wise exposure, the row-kernel and column-kernel convolutions were leveraged to extract attributes along horizontal and vertical axes. RSC-Net [100] improved URS-CNN [27] from 2 degrees of freedom (DoF) to 6-DoF and presents a structure-and-motion-aware RS correction model, where the camera scanline velocity and depth were estimated. Compared to URS-CNN [27], RSC-Net [100] further reasoned about the concealed motion between the scanlines as well as the scene structure as shown in Figure 6. To bridge the spatiotemporal connection between RS and GS, EvUnroll [42] exploited the neuromorphic events to correct the RS effect. Event cameras can overcome a number of drawbacks of conventional frame-based activities for dynamic situations with quick motion due to their high temporal resolution property with microsecond-level sensitivity.
4.2.2 Multi-frame-based Solution
Most multi-frame-based solutions are based on the reconstruction paradigm, they mainly devote to contributing how to represent the dense displacement field between RS and global GS images and accurately warp the RS domain to the GS domain. For the first time, DeepUnrollNet [49] proposed an end-to-end network for two consecutive rolling shutter images using a differentiable forward warping module. In this method, a motion estimation network is used to estimate the dense displacement field from a rolling shutter image to its matching global shutter image. The second contribution of DeepUnrollNet [49] is to construct two novel datasets: the Fastec-RS dataset and the Carla-RS dataset. Furthermore, JCD [153] jointly considered the rolling shutter correction and deblurring (RSCD) techniques, which largely exist in the medium and long exposure cases of rolling shutter cameras. It applied bi-directional warping streams to compensate for the displacement while keeping the non-warped deblurring stream to restore details. The authors also contributed a real-world dataset using a well-designed beam-splitter acquisition system, BS-RSCD, which includes both ego-motion and object motion in dynamic scenes. SUNet [159] extended DeepUnrollNet [49] from the middle time of the second frame () into the intermediate time of two frames (). By using PWC-Net [205], SUNet [159] estimated the symmetric undistortion fields and reconstructed the potential GS frames by a time-centered GS image decoder network. To effectively reduce the misalignment between the contexts warped from two consecutive RS images, the context-aware undistortion flow estimator and the symmetric consistency enforcement were designed. To achieve a higher frame rate, Fan et al. [158] generated a GS video from two consecutive RS images based on the scanline-dependent nature of the RS camera. In particular, they first analyzed the inherent connection between bidirectional RS undistortion flow and optical flow, demonstrating the RS undistortion flow map has a more pronounced scanline dependency than the isotropically smooth optical flow map. Then, they developed the bidirectional undistortion flows to describe the pixel-wise RS-aware displacement, and further devised a computation technique for the mutual conversion between different RS undistortion flows corresponding to various scanlines. To eliminate the inaccurate displacement field estimation and error-prone warping problems in previous methods, AW-RSC [166] proposed to predict multiple fields and adaptively warped the learned RS features into global shutter counterparts. Using a coarse-to-fine approach, these warped features were combined and generated to precise global shutter frames as shown in Figure 7. Compared to previous works [49, 153, 159, 158], the warping operation consisting of adaptive multi-head attention and a convolutional block in AW-RSC [166] is learnable and effective. In addition, AW-RSC [166] contributed a real-world rolling shutter correction dataset: BS-RSC, where the RS videos with corresponding GS ground truth are captured simultaneously with a beam-splitter-based acquisition system.
4.3 Discussion
4.3.1 Technique Summary
The deep learning works on wide-angle camera and roll shutter calibration share a similar technique pipeline. Along this research trend, most early literature begins with the regression-based solution [26, 40, 27]. The subsequent works innovated the traditional calibration with a reconstruction perspective [34, 35, 121, 49], which directly learns the displacement field to rectify the uncalibrated input. For higher accuracy of calibration, a more intuitive displacement field, and more effective warping strategy have been developed [145, 166, 153, 158]. To fit the distribution of different distortions, some works designed different shapes of the convolutional kernel [27] or transformed the convolved coordinates [143].
Existing works devoted themselves to designing more powerful networks and introducing more diverse features to facilitate calibration performance. Increasingly more methods focused on the geometry priors of the distortion [121, 119, 81]. These priors can be directly weighted into the convolutional layers or used to supervise network training, promoting the learning model to converge faster.
4.3.2 Future Effort
(1) The development of wide-angle camera calibration and roll shutter camera calibration can promote each other. For instance, the well-studied multi-frame-based solution in roll shutter calibration is able to inspire wide-angle calibration. The same object located at different sequences could provide useful priors regarding to radial distortion. Additionally, the elaborate studies of the displacement field and warping layer [166, 153, 158] have the potential to motivate the development of wide-angle camera calibration and other fields. Furthermore, the investigation of geometric priors in wide-angle calibration could also improve the interpretability of the network in roll shutter calibration.
(2) Most methods synthesize their training dataset based on random samples from all camera parameters. However, for the images captured by real lenses, the distribution of camera parameters probably locates at a potential manifold [88]. Learning on a label-redundant calibration dataset makes the training process inefficient. Thus, exploring a practical sampling strategy for the synthesized dataset could be a meaningful task in the future direction.
(3) To overcome the ill-posed problem of single-frame calibration, introducing other high-precision sensors can compensate for the current calibration performance, such as event cameras [42]. With the rapid development of vision sensors, joint calibration using multiple sensors is valuable. Consequently, more cross-modal and multi-modal fusion techniques will be investigated along this research way.
5 Cross-View Model
The existing deep calibration methods can estimate the specific camera parameters from a single camera. In fact, there can be more complicated parameter representations in multi-camera circumstances. For example, in the multi-view model, the fundamental matrix and essential matrix describe the epipolar geometry and they are intricately tangled with intrinsics and extrinsics. The homography depicts the pixel-level correspondences between different views. In addition to intrinsics and extrinsics, it is also intertwined with depth. Among these complex parameter representations, homography is the most widely leveraged in practical applications and its related learning-based methods are the most investigated. To this end, we mainly focus on the review of deep homography estimation solutions for the cross-view model and they can be divided into three categories: direct, cascaded, and iterative solution.
5.1 Direct Solution
We review the direct deep homography solutions from the perspective of different parameterizations, including the classical 4-pt parameterization and other parameterizations.
5.1.1 4-pt Parameterization
Deep homography estimation is first proposed in DHN[28], where a VGG-style network is adopted to predict the 4-pt parameterization . To train and evaluate the network, a synthetic dataset named Warped MS-COCO is created to provide ground truth 4-pt parameterization . The pipeline is illustrated in Fig. 8(a), and the objective function is formulated as :
Then the 4-pt parameterization can be solved as a homography matrix using normalized DLT[206]. However, DHN is limited to synthetic datasets where the ground truth can be generated for free or requires costly labeling of real-world datasets. Subsequently, the first unsupervised solution named UDHN[52] is proposed to address this problem. As shown in Fig. 8(c), it used the same network architecture as DHN and defined an unsupervised loss function by minimizing the average photometric error motivated by traditional methods[207]:
where and denote the operations of warping via homography parameters and extracting an image patch, respectively. and are the original images with overlapping regions. The input of UDHN is a pair of image patches, but it warps the original images when calculating the loss. In this manner, it avoids the adverse effects of invalid pixels after warping and lifts the magnitude of pixel supervision. To gain accuracy and speed with a tiny model, Chen et al. proposed ShuffleHomoNet [148], which integrates ShuffleNet compressed units[208] and location-aware pooling[84] into a lightweight model. To further handle large displacement, a multi-scale weight-sharing version is exploited by extracting multi-scale feature representations and adaptively fusing multi-scale predictions. However, the homography cannot perfectly align images with parallax caused by non-planar structures with non-overlapping camera centers. To deal with parallax, CA-UDHN[53] designs learnable attention masks to overlook the parallax regions, contributing to better background plane alignment. Besides, the 4-pt homography can be extended to meshflow[56] to realize non-planar accurate alignment.
5.1.2 Other Parameterizations
In addition to 4-pt parameterization, the homography can be parameterized as other formulations. To better utilize homography invertibility, Wang et al. proposed SSR-Net [58]. They established the invertibility constraint through a conventional matrix representation in a cyclic manner. Zeng et al. [85] argued that the 4-point parameterization regressed by a fully-connected layer can harm the spatial order of the corners and be susceptible to perturbations, since four points are the minimum requirement to solve the homography. To address these issues, they formulated the parameterization as a perspective field (PF) that models pixel-to-pixel bijection and designed a PFNet. This extends the displacements of the four vertices to as many dense pixel points as possible. The homography can then be solved using RANSAC [209] with outlier filtering, enabling robust estimation by utilizing dense correspondences. Nevertheless, dense correspondences lead to a significant increase in the computational complexity of RANSAC. Furthermore, Ye et al.[54] proposed an 8-DOF flow representation without extra post-processing, which has a size of in an 8D subspace constrained by the homography. To represent arbitrary homography flows in this subspace, 8 flow bases are defined, and the proposed BasesHomo is to predict the coefficients for the flow bases. To obtain desirable bases, BasesHomo first generates 8 homography flows by modifying every single entry of an identity homography matrix except for the last entry. Then, these flows are normalized by their largest flow magnitude followed by a QR decomposition, enforcing all the bases normalized and orthogonal.
5.2 Cascaded Solution
Direct solutions explore various homography parameterizations with simple network structures, while the cascaded ones focus on complex designs of network architectures.
In HierarchicalNet[72], Nowruzi et al. hold that the warped images can be regarded as the input of another network. Therefore they stacked the networks sequentially to reduce the error bounds of the estimate. Based on HierarchicalNet, SRHEN [136] introduced the cost volume[205] to the cascaded network, measuring the feature correlation by cosine distance and formulating it as a volume. The stacked networks and cost volume increase the performance, but they cannot handle the dynamic scenes. MHN [113] developed a multi-scale neural network and proposed to learn homography estimation and dynamic content detection simultaneously. Moreover, to tackle the cross-resolution problem, LocalTrans [147] formulated it as a multimodal problem and proposed a local transformer network embedded within a multiscale structure to explicitly learn correspondences between the multimodal inputs. These inputs include images with different resolutions, and LocalTrans achieved superior performance on cross-resolution cases with a resolution gap of up to 10x. All the solutions mentioned above leverage image pyramids to progressively enhance the ability to address large displacements. However, every image pair at each level requires a unique feature extraction network, resulting in the redundancy of feature maps. To alleviate this problem, some researchers[210, 149, 44, 55] replaced image pyramids with feature pyramids. Specifically, they warped the feature maps directly instead of images to avoid excessive feature extraction networks. To address the low-overlap homography estimation problem in real-world images[149], Nie et al.[149] modified the unsupervised constraint (Eq. 3) to adapt to low-overlap scenes:
where is an all-one matrix with the same size as or . It solved the low-overlap problem by taking the original images as network input and ablating the corresponding pixels of to the invalid pixels of warped . To solve the non-planar homography estimation problem, DAMG-Homo[44] proposed backward multi-gird deformation with contextual correlation to align parallax images. Compared with traditional cost volume, the proposed contextual correlation helped to reach better accuracy with lower computational complexity. Another way to address the non-planar problem is to focus on the dominant plane. In HomoGAN [55], an unsupervised GAN is proposed to impose a coplanarity constraint on the predicted homography, as shown in Figure 9. To implement this approach, a generator is used to predict masks of aligned regions, while a discriminator is used to determine whether two masked feature maps were produced by a single homography.
5.3 Iterative Solution
Compared with cascaded methods, iterative solutions achieve higher accuracy by iteratively optimizing the last estimation. Lucas-Kanade (LK) algorithm[207] is usually used in image registration to estimate the parameterized warps iteratively, such as affine transformation, optical flow, etc. It aims at the incremental update of warp parameters every iteration by minimizing the sum of squared error between a template image and an input image :
However, when optimizing Eq. 5 using first-order Taylor expansion, should be recomputed every iteration because varies with . To avoid this problem, the inverse compositional (IC) LK algorithm[211], an equivalence to LK algorithm, can be used to reformulate the optimization goal as follows:
After linearizing Eq. 6 with first-order Taylor expansion, we compute instead of , which would not vary every iteration.
To combine the advantages of deep learning with IC-LK iterator, CLKN [71] conducted LK iterative optimization on semantic feature maps extracted by CNNs as follows:
where and are the feature maps of the template and input images. Then, they enforced the network to run a single iteration with a hinge loss, while the network runs multiple iterations until the stopping condition is met in the testing stage. Besides, CLKN stacked three similar LK networks to further boost the performance by treating the output of the last LK network as the initial warp parameters of the next LK network. From Eq. 7, the IC-LK algorithm heavily relied on feature maps, which tend to fail in multimodal images. Instead, DLKFM [146] constructed a single-channel feature map by using the eigenvalues of the local covariance matrix on the output tensor. To learn DLKFM, it designed two special constraint terms to align multimodal feature maps and contribute to convergence.
However, LK-based algorithms can fail if the Jacobian matrix is rank-deficient [212]. Additionally, the IC-LK iterator is untrainable, which means this drawback is theoretically unavoidable. To address this issue, a completely trainable iterative homography network (IHN) [165] was proposed. Inspired by RAFT [213], IHN updates the cost volume to refine the estimated homography using the same estimator repeatedly every iteration. Furthermore, IHN can handle dynamic scenes by producing an inlier mask in the estimator without requiring extra supervision.
5.4 Discussion
5.4.1 Technique Summary
The above works are devoted to exploring different homography parameterizations such as 4-pt parameterization[28], perspective field[85], and motion bases representation[54], which contributes to better convergence and performance. Other works tend to design various network architectures. In particular, cascaded and iterative solutions are proposed to refine the performance progressively, which can be further combined together to reach higher accuracy. To make the methods more practical, various challenging problems are preliminarily addressed, such as cross resolutions[147], multiple modalities[146, 165], dynamic objects[113, 165], and non-planar scenes[53, 55, 44], etc.
5.4.2 Challenge and Future Effort
We summarize the existing challenges as follows:
(1) Many homography estimation solutions are designed for fixed resolutions, while real-world applications often involve much more flexible resolutions. When pre-trained models are applied to images with different resolutions, performance can dramatically drop due to the need for input resizing to satisfy the regulated resolution.
(2) Unlike optical flow estimation, which assumes small motions between images, homography estimation often deals with images that have significantly low-overlap rates. In such cases, existing methods may exhibit inferior performance due to limited receptive fields.
(3) Existing methods address the parallax or dynamic objects by learning to reject outliers in the feature extractor[53], cost volume[214], or estimator[165]. However, it is still unclear which stage is more appropriate for outlier rejection.
Based on the challenges we have discussed, some potential research directions for future efforts can be identified:
(1) To overcome the first challenge, we can design various strategies to enhance resolution robustness, such as resolution-related data augmentation, and continual learning on multiple datasets with different resolutions. Besides, we can also formulate a resolution-free parameterization form. The perspective field [85] is a typical case, which represents the homography as dense correspondences with the same resolution as input images. But it requires RANSAC as the post-processing approach, introducing extra computational complexity, especially in the case of extensive correspondences. Therefore, a resolution-free and efficient parameterization form should be explored.
(2) To enhance the performance in low-overlap rate, the main insight is to increase the receptive fields of a network. To this end, the cross-attention module of the transformer explicitly leverages the long-range correlation to eliminate short-range inductive bias[215]. On the other hand, we can exploit beneficial varieties of cost volume to integrate feature correlation [44, 165].
(3) As there is no interaction between different image features in the feature extractor, it is reasonable to assume that outlier rejection should occur after feature extraction. It is not possible to identify outliers within a single image as the depth alone cannot be used as an outlier cue. For example, images captured by purely rotated cameras do not contain parallax outliers. Additionally, it seems intuitive to learn the capability of outlier rejection by combining global and local correlation, similar to the insight of RANSAC.
6 Cross-Sensor Model
Multi-sensor calibration estimates intrinsic and extrinsic parameters of multiple sensors like cameras, LiDARs, and IMUs. This ensures that data from different sensors are synchronized and registered in a common coordinate system, allowing them to be fused together for a more accurate representation of the environment. Accurate multi-sensor calibration is crucial for applications like autonomous driving and robotics, where reliable sensor fusion is necessary for safe and efficient operation.
In this part, we mainly review the literature on learning-based camera-LiDAR calibration, i.e., predicting the 6-DoF rigid body transformation between a camera and a 3D LiDAR, without requiring any presence of specific features or landmarks in the implementation. Like calibration works on other types of cameras/systems, this research field can also be classified into regression-based solutions and flow/reconstruction-based solutions. But we are prone to follow the special matching principle in camera-LiDAR calibration and divide the existing learning-based literature into three categories: pixel-level solution, semantics-level solution, and object/keypoint-level solution.
6.1 Pixel-level Solution
The first deep learning technique in camera-LiDAR calibration, RegNet [30], used CNNs to combine feature extraction, feature matching, and global regression to infer the 6-DoF extrinsic parameters. It processed the RGB and LiDAR depth map separately and branched two parallel data network streams. Then, a specific correlation layer was proposed to convolve the stacked LiDAR and RGB features as a joint representation. After this feature matching, the global information fusion and parameter regression were achieved by two fully connected layers with a Euclidean loss function. Motivated by this work, the subsequent works made a further step into more accurate camera-LiDAR calibration in terms of the geometric constraint [86, 131], temporal correlation [131], loss design [130], feature extraction [182], feature matching [155, 135], feature fusion [182], and calibration representation [156, 179].
For example, as shown in Figure 10, CalibNet [86] designed a network to predict calibration parameters that maximize the geometric and photometric consistency of images and point clouds, solving the underlying physical problem by 3D Spatial Transformers [216]. To refine the calibration model, CalibRCNN [131] presented a synthetic view and an epipolar geometry constraint to measure the photometric and geometric inaccuracies between consecutive frames, of which the temporal information learned by the LSTM network has been investigated in the learning-based camera-LiDAR calibration for the first time. Since the output space of the LiDAR-camera calibration is on the 3D Special Euclidean Group () rather than the normal Euclidean space, RGGNet [130] considered Riemannian geometry constraints in the loss function, namely, used a geodesic distance equipped with left-invariant Riemannian metrics to optimize the calibration network. LCCNet [155] exploited the cost volume layer to learn the correlation between the image and the depth transformed by the point cloud. Because the depth map ignores the 3D geometric structure of the point cloud, FusionNet [182] leveraged PointNet++ [217] to directly learn the features from the 3D point cloud. Subsequently, a feature fusion with Ball Query [217] and attention strategy was proposed to effectively fuse the features of images and point clouds.
CFNet [156] first proposed the calibration flow for camera-LiDAR calibration, which represents the deviation between the positions of initial projected 2D points and ground truth. Compared to directly predicting extrinsic parameters, learning the calibration flow helped the network to understand the underlying geometric constraint. To build precise 2D-3D correspondences, CFNet [156] corrected the originally projected points using the estimated calibration flow. Then the efficient Perspective-n-Point (EPnP) algorithm was applied to calculate the final extrinsic parameters by RANSAC. Because RANSAC is nondifferentiable, DXQ-Net [179] further presented a probabilistic model for LiDAR-camera calibration flow, which estimates the uncertainty to measure the quality of LiDAR-camera data association. Then, the differentiable pose estimation module was designed for solving extrinsic parameters, back-propagating the extrinsic error to the flow prediction network.
6.2 Semantics-level Solution
Semantic features can be well learned and represented by deep neural networks. A perfect calibration enables to accurately align the same instance in different sensors. To this end, some works [134, 132, 160, 45] explored to guide the camera-LiDAR calibration with the semantic information. SOIC [134] calibrated and transforms the initialization issue into the semantic centroids’ PnP problem using semantic information. Since the 3D semantic centroids of the point cloud and the 2D semantic centroids of the picture cannot match precisely, a matching constraint cost function based on the semantic components was presented. SSI-Calib [132] reformulated the calibration as an optimization problem with a novel calibration quality metric based on semantic features. Then, a non-monotonic subgradient ascent algorithm was proposed to calculate the calibration parameters. Other works utilized the off-the-shelf segmentation networks for point cloud and image, and optimized the calibration parameters by minimizing semantic alignment loss in single-direction [160] and bi-direction [45].
6.3 Object/Keypoint-level Solution
ATOP [181] designed an attention-based object-level matching network, i.e., Cross-Modal Matching Network to explore the overlapped FoV between camera and LiDAR, which facilitated generating the 2D-3D object-level correspondences. 2D and 3D object proposals were detected by YOLOv4 [218] and PointPillar [219]. Then, two cascaded PSO-based algorithms [220] were devised to estimate the calibration extrinsic parameters in the optimization stage. Using the deep declarative network (DDN) [221], RGKCNet [183] combined the standard neural layer and a PnP solver in the same network, formulating the 2D–3D data association and pose estimation as a bilevel optimization problem. Therefore, both the feature extraction capability of the convolutional layer and the conventional geometric solver can be employed. Microsoft’s human keypoint extraction network [222] was applied to detect the 2D–3D matching keypoints. Additionally, RGKCNet [183] presented a learnable weight layer that determines the keypoints involved in the solver, enabling the whole pipeline to be trained end-to-end.
6.4 Discussion
6.4.1 Technique Summary
The current method can be briefly classified based on the principle of building 2D and 3D matching, namely, the calibration target. In summary, most pixel-level solutions utilized the end-to-end framework to address this task. While these solutions delivered satisfactory performances on specific datasets, their generalization abilities are limited. Semantics-level and object/keypoint-level methods derived from traditional calibration offered both acceptable performances and generalization abilities. However, they heavily relied on the quality of fore-end feature extraction.
6.4.2 Research Trend
(1) Network architecture is becoming more complex with the use of different structures for feature extraction, matching, and fusion. Current methods employ strategies like multi-scale feature extraction, cross-modal interaction, cost-volume establishment, and confidence-guided fusion.
(2) Directly regressing 6-DoF parameters yields weak generalization ability. To overcome this, intermediate representations like calibration flow have been introduced. Additionally, calibration flow can handle non-rigid transformations that are common in real-world applications.
(3) Traditional methods require specific environments but have well-designed strategies. To balance accuracy and generalization, a combination of geometric solving algorithms and learning methods has been investigated.
6.4.3 Future Effort
(1) Camera-LiDAR calibration methods typically rely on datasets like KITTI, which provide only initial extrinsic parameters. To create a decalibration dataset, researchers add noise transformations to the initial extrinsics, but this approach assumes a fixed position camera-LiDAR system with miscalibration. In real-world applications, the camera-LiDAR relative pose varies, making it challenging to collect large-scale real data with ground truth extrinsics. To address this challenge, generating synthetic camera-LiDAR data using simulation systems could be a valuable solution.
(2) To optimize the combination of networks and traditional solutions, a more compact approach is needed. Current methods mainly use networks as feature extractors, resulting in non-end-to-end pipelines with inadequate feature extraction adjustments for calibration. A deep declarative network (DDN) is a promising framework for making the entire pipeline differentiable. The aggregation of learning and traditional methods can be optimized using DDN.
(3) The most important aspect of camera-LiDAR calibration is 2D-3D matching. To achieve this, the point cloud is commonly transformed into a depth image. However, large deviations in extrinsic simulation can result in detail loss. With the great development of Transformer and cross-modal techniques, we believe leveraging Transformer to directly learn the features of image and point cloud in the same pipeline could facilitate better 2D-3D matching.
7 Benchmark
As there is no public and unified benchmark in learning-based camera calibration, we contribute a dataset that can serve as a platform for generalization evaluations. In this dataset, the images and videos are captured by different cameras under diverse scenes, including simulation environments and real-world settings. Additionally, we provide the calibration ground truth, parameter label, and visual clues in this dataset based on different conditions. Figure 11 shows some samples of our collected dataset. Please refer to the evaluation of representative calibration methods on this benchmark in supplementary material.
Standard Model. We collected 300 high-resolution images on the Internet, captured by popular digital cameras such as Canon, Fujifilm, Nikon, Olympus, Sigma, Sony, etc. For each image, we provide the specific focal length of its lens. We have included a diverse range of subjects, including landscapes, portraits, wildlife, architecture, etc. The range of focal length is from 4.5mm to 600mm.
Distortion Model. We created a comprehensive dataset for the distortion camera model, with a focus on wide-angle cameras. The dataset is comprised of three subcategories. The first is a synthetic dataset, which was generated using the widely-used 4th order polynomial model. It contains both circular and rectangular structures, with 1,000 distortion-rectification image pairs. The second subcategory consists of data captured under real-world settings, derived from the raw calibration data for around 40 types of wide-angle cameras. For each calibration data, the intrinsics, extrinsics, and distortion coefficients are available. Finally, we exploit a car equipped with different cameras to capture video sequences. The scenes cover both indoor and outdoor environments, including daytime and nighttime footage.
Cross-View Model. We selected 500 testing samples at random from each of four representative datasets (MS-COCO [28], GoogleEarch [146], GoogleMap [146], CAHomo [53]) to create a dataset for the cross-view model. It covers a range of scenarios: MS-COCO provides natural synthetic data, GoogleEarch contains aerial synthetic data, and GoogleMap offers multi-modal synthetic data. Parallax is not a factor in these three datasets, while CAHomo provides real-world data with non-planar scenes. To standardize the dataset, we converted all images to a unified format and recorded the matched points between two views. In MS-COCO, GoogleEarch, and GoogleMap, we used four vertices of the images as the matched points. In CAHomo, we identified six matched key points within the same plane.
Cross-Sensor Model. We collected RGB and point cloud data from Apollo [223], DAIR-V2X [224], KITTI [77], KUCL [225], NuScenes [226], and ONCE [227]. Around 300 data pairs with calibration parameters are included in each category. The datasets are captured in different countries to provide enough variety. Each dataset has a different sensor setup, obtaining camera-LiDAR data with varying image resolution, LiDAR scan pattern, and camera-LiDAR relative location. The image resolution ranges from 24482048 to 1242375, while the LiDAR sensors are from Velodyne and Hesai, with 16, 32, 40, 64, and 128 beams. They include not only normal surrounding multi-view images but also small baseline multi-view data. Additionally, we also added random disturbance of around 20 degrees rotation and 1.5 meters translation based on classical settings [30] to simulate vibration and collision.
8 Future Research Directions
Camera calibration is a fundamental and challenging research topic. From the above technical reviews and limitation analysis, we can conclude there is still room for improvement with deep learning. From Section 3 to Section 6, specific future efforts are discussed for each model. In this section, we suggest more general future research directions.
8.1 Sequences
Bundle adjustment is a well-established technique central to Multi-View Stereo (MVS) and Simultaneous Localization and Mapping (SLAM) using multi-view constraints. Traditional bundle adjustment focuses on pose estimation, often under the assumption of pre-calibrated cameras, thus sidelining the nuances of camera parameter fine-tuning. While learning-based camera calibration has made significant strides, most methods are tailored for a single image. We have highlighted intrinsics calibration to underscore how sequence constraints bolster prediction accuracy. Notably, there is a burgeoning interest in integrating bundle adjustment into end-to-end deep learning pipelines. By transitioning from conventional keypoint extraction and matching to learning-based methods, recent works [228, 229, 230, 231, 232] propose differentiable bundle adjustment layers to refine pose, depth, and camera parameters together. Consequently, there is immense potential in further harnessing sequence constraints for accurate calibration. Current methods combine front-end matching with a back-end solver, which can be inefficient and unreliable in cases like fast motion. We suggest separating front-end and back-end refinements, using large models for features and introducing more trainable parameters in optimization.
8.2 Learning Target
Due to the implicit relationship to image features, conventional calibration objectives can be challenging for neural networks to learn. To this end, some works have developed novel learning targets that replace conventional calibration objectives, providing learning-friendly representations for neural networks. Additionally, intermediate geometric representations have been presented to bridge the gap between image features and calibration objectives, such as reflective amplitude coefficient maps [128], rectification flow [37], surface geometry [89], and normal flow [170], etc. Looking ahead to the future development of this community, we believe there is still great potential for designing more explicit and reasonable learning targets for calibration objectives.
8.3 Pre-training
Pre-training on ImageNet [69] has become a widely used strategy in deep learning. However, recent studies [96] have shown that this approach provides less benefit for specific camera calibration tasks, such as wide-angle camera calibration. This is due to two main reasons: the data gap and the task gap. The ImageNet dataset only contains perspective images without distortions, making the initialized weights of neural networks irrelevant to distortion models. Furthermore, He et al. [233] demonstrated that the task of ImageNet pre-training has limited benefits when the final task is more sensitive to localization. As a result, the performance of extrinsics estimation may be impacted by this task gap. Moreover, pre-training beyond a single image and a single modality, to our knowledge, has not been thoroughly investigated in the related field. We suggest that designing a customized pre-training strategy for learning-based camera calibration is an interesting area of research.
8.4 Implicit Unified Model
Deep learning-based camera calibration methods use traditional parametric camera models, which lack the flexibility to fit complex situations. Non-parametric camera models relate each pixel to its corresponding 3D observation ray, overcoming parametric model limitations. However, they require strict calibration targets and are more complex for undistortion, projection, and unprojection. Deep learning methods show potential for calibration tasks, making non-parametric models worth revisiting and potentially replacing parametric models in the future. Moreover, they allow for implicit and unified calibration, fitting all camera types through pixel-level regression and avoiding explicit feature extraction and geometry solving. Researchers combined the advantages of implicit and unified representation with the Neural Radiance Field (NeRF) for reconstructing 3D structures and synthesizing novel views. Self-calibration NeRF [191] has been proposed for generic cameras with arbitrary non-linear distortions, and end-to-end pipelines have been explored to learn depth and ego-motion without calibration targets. We believe the implicit and unified camera models could be used to optimize learning-based algorithms or integrated into downstream 3D vision tasks.
9 Conclusion
In this paper, we present a comprehensive survey of the recent efforts in the area of deep learning-based camera calibration. Our survey covers conventional camera models, classified learning paradigms and learning strategies, detailed reviews of the state-of-the-art approach, a public benchmark, and future research directions. To exhibit the development process and link the connections between existing works, we provide a fine-grained taxonomy that categorizes literature by jointly considering camera models and applications. Moreover, the relationships, strengths, distinctions, and limitations are thoroughly discussed in each category. An open-source repository will keep updating regularly with new works and datasets. We hope that this survey could promote future research in this field.
Acknowledgment
We thank Leidong Qin and Shangrong Yang at Beijing Jiaotong University for the partial dataset collection. We thank Jinlong Fan at the University of Sydney for the insightful discussion.